name direction vertex Axis of symmetry focus diectrix of the parabola
x-1=1/4(y-8)^2
please help

Answer by lwsshak3(11628)

name direction vertex, Axis of symmetry, focus, directrix of the parabola
x-1=1/4(y-8)^2
(y-8)^2=4(x-1)
This is an equation of a parabola that opens rightward.
Its basic form: (y-k)^2=4p(x-k), (h,k)=coordinates of the vertex
For given parabola:
vertex: (1,8)
axis of symmetry: y=8
4p=4
p=1
focus(2,8)
directrix: x=0, or y-axis
see graph below:
y=±(4x-4)^.5+8